• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 18:12:53 on July 02, 2010

    Entry number 329476

    keyword=Suggested VDC checkout

    Suggested VDC checks for tonight's detector checkout:
    
    For both arms, set VDC HV to -3500 V, thresholds to 2.0 V before data
    taking.
    
    Take calibration data in both arms with white spectrum (e.g. 12C target)
    We need at least 1M singles triggers (the S2m triggers) in each arm.
    
    (1) Verify L-HRS VDCs still good after timing changes (10 min)
    
    Check time offset in raw TDC spectra ("L-arm VDC TDC" in onlineGUI).
    Right edge of spectrum (right side of big peak) should be around 1220,
    and the spectrum must not be cut off. If the offset is far off, email me
    to do a recalibration (or do one yourself - see halog 328603).
    
    Also, check wiremaps and efficiencies for any obvious problems that
    may have crept in (disconnected cables etc.)
    
    (2) Check R-HRS VDC (10 min)
    
    (a) Check wiremaps ("R-arm VDC wires"). Must be smooth, no peaks, edges,
    sawtooth patterns, dead or hot channels or channel groups. 
    
    (b) Check TDC spectra, determine offset. R-HRS database has not been
    updated yet. Time offsets in the database are around 1685. Perform TDC
    calibration to adjust.
    
    (c) Check R-HRS VDC efficiency plots (R-arm VDC efficiency (zoomed).
    Efficiencies should be 99% or better everywhere.
    
    Any problems discovered probabyl require access to replace bad
    front-end cards, fix cable problems etc.
    
    Once VDCs in both arms are working satisfactorily, proceed:
    
    (3) HV and threshold scans (1-2 hrs)
    
    (a) For each of the follwoing HV and threshold settings, take a 1-2 
    minute R-arm run at low rates (<250 kHz) and check the R-HRS VDC
    efficiency:
    
    HV (V)   thresh (V)
    -3450     2.0
    -3400     2.0
     efficiency should drop visibly, if not, skip next 2 pts and do
     do -3350 and -3300 instead
    -3475     2.0
    -3425     2.0
    -3500     3.0
    -3500     2.5
    -3450     2.5
    -3450     3.0
    
    Return to HV=-3500 V and thresh=2.0 V in both arms, unless the tests
    suggest a better setting for the R-HRS.
    
    (b) High-rate checks. Estabish high-rates in both arms (preferably
    >1 MHz singles in R-HRS, 5-6 MHz in L-HRS) as conditions permit. 
    Adjust prescales for <20% deadtime. Take 1-2 minute runs with both arms
    (goal >=100k events in each arm) for each of the following VDC HV and 
    threshold settings:
    
    HV (V)      thresh (V)
    -3500       3.5
    -3500       3.0
    -3500       2.5
    -3500       2.0
    -3500       1.5
    -3450       1.5
    -3450       2.0
    -3450       2.5
    -3450       3.0
    -3450       3.5
    
    Return to HV=-3500V and threshold=2.0V at end of checks.
    
    


    A copy of this log entry has been emailed to: riordan, bogdanw, ntoro@stanford.edu, schuster@slac.stanford.edu